Write an equation in point slope form for the perpendicular bisector of a segment with endpoints M(-5,4), N(1,-2)
Answer by mananth(16946)

M(-5,4), N(1,-2)
Mid point of MN= (-5+1)/2 , (4-2)/2
=(-2,1)
The perpendicular bisector passes through this point and it is perpendicular to the line joining MN
slope of MN
(y2-y1)/(x2-x1)
(-2-4)/(1-(-5)
-6/6
=-1
slope of perpendicular bisector is +1
The line has slope of 1 and passes through (-2,1)
y-y1=m(x-x1)
(y-1)=1(x-(-2))
y-1=x+2
y=x+3